OpenStudy (anonymous):

-3 4/5+ (-1 7/10) write the sum as a mixed number? explain.

OpenStudy (anonymous):

i got -3 11/15 but i need help explaining how to change it to a mixed number

OpenStudy (anonymous):

it says to convert the mixed numbers to fractions

OpenStudy (anonymous):

then find the common denominator and add the fraction

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

\[-3 \frac{ 4 }{5 }+(-1 \frac{ 7 }{ 10 })= \frac{ -19 }{ 5 }- \frac{ 17 }{ 10 }=\frac{ -38 }{ 10 }-\frac{ 17 }{ 10 }=\frac{ -38-17 }{ 10 }=\frac{ -55 }{ 10 }=-5\frac{ 5 }{ 10 }=-5\frac{ 1 }{ 2 }\]

TheSmartOne (thesmartone):

\[=-5\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

im confused how did you get -19

OpenStudy (anonymous):

sorry i dont have a calculator with me

OpenStudy (anonymous):

it said it was 2 1/10?

TheSmartOne (thesmartone):

Oh..

OpenStudy (anonymous):

lol positive 19 and -17 you had it set up right just the positive and negatives were wrong

OpenStudy (anonymous):

then you had to go (19x2 + -17/10) = (38/10+-17/10)= 38+ -17/10) and ended up being 2 1/10 5x2

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

4/7 -(-2/7) find the difference. so would it be 8/7 or 8/49?

TheSmartOne (thesmartone):

it would be 4/7+2/7=6/7

TheSmartOne (thesmartone):

http://www.helpwithfractions.com/fraction-calculator/

OpenStudy (anonymous):

ohh okay and thanks :)

OpenStudy (anonymous):

i already have that link lol

OpenStudy (anonymous):

i put my problem in there and said it was wrong

OpenStudy (anonymous):

its 2 5/8 - (-1 1/4) and i got 3 3/8 so do i simplify? cause all it says is to put it as a mixed number?

TheSmartOne (thesmartone):

it is simplified...

OpenStudy (anonymous):

somehow i got it wrong

OpenStudy (anonymous):

my online class is stupid

OpenStudy (anonymous):

it said its 3 7/8?